Recent Log Entries

5th Pass Horizontal Separation Checked -- 17.3mm!

Log Number: 3626749Logbook(s): Entry Makers: moserBody: 

Figure 1: C-beam. Figure 2: B-beam. Figure 3: D-beam. Figure 4: All 3 beams.

Figure 5: C_beam on BPMs. Going to move this up to -11. Figure 6: Adjustment to get to -11mm.

Figure 7: B-beam. Figure 8: D-beam. Holy separators we are over-separated here too! -11mm to 6.3 = 17.3mm of separation! Awesome!

Going to live with this for now. dogleg calculator

Figures 9 and 10: 5th Pass horizontal separator.

FSD Masking Tool: Successful System Masking for: BeamLossMonitor

Log Number: 3626742Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: moser
Areas: All

BeamLossMonitor Devices:
ILM1L02
ILM0I07
ILM6R08
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0R04
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R06
ILM0R08
ILM0R09
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T09
ILM4C01
ILM4C02
ILM4C03
ILM1C01
ILM1C05<

2nd Pass Horizontal Separation Good, Perhaps Even Too Good!

Log Number: 3626735Logbook(s): Entry Makers: moserBody: 

Figure 1: 4E02 viewer with A-beam only. Figure 2: +C. Figure 3: A+B. Figure 4: A+D. Figure 5: A+B+C+D.

Figure 6: C-beam. Figure 7: A-beam. If anything, we have too much separation -> 17mm rather than 16.5mm. Can live with this. Figure 8: B-beam. Figure 9: D-beam.

If we find we are getting 4E02 BLM trips we'll reduce the separation to design 16.5mm.

Figure 10 and 11: 2nd Pass RF Separator.

Injector restore after spot move

Log Number: 3626732Logbook(s): Entry Makers: kazimihoflerBody: 

After spot move, we steered to minimize interception on A3.

Figs. 1-4 Interception for beams C (85 uA), A (37 uA), B (14 uA), and D (25 uA), respectively.
Fig. 5 C orbit
Figs. 6-9 A, B, C, and D orbits relative to C

allsave 19455 with zero pos using C beam

FSD Masking Tool: Successful System Masking for: BeamLossMonitor

Log Number: 3626731Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: lrichard
Areas: All

BeamLossMonitor Devices:
ILM1L02
ILM0I07
ILM6R08
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0R04
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R06
ILM0R08
ILM0R09
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T09
ILM4C01
ILM4C02
ILM4C03
ILM1C01
ILM1C

Autolog save

Log Number: 3626723Logbook(s): Tags: Body: saver ran from opsl04

Saved: kazimi 11-15-18 15:36:41 {A:2 B:5 C:5 E:1050 A=BS B=BS C=CW D=BS 0R08=0.00 0R08 dump in. A B C dump(s) in.

CREW CHIEF SIGN IN

Log Number: 3626722Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 11-15-18
Time : 15:29:56
Day : Thursday
Shift : Swing

Crew Chief : Comer
Program Deputy: Vasilauskis
Team : F
Arm : Turner
Designated SSO: Cade
LERF Operator:
Operators : Cade, Turner, Garcia

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

LEM North Linac

Log Number: 3626721Logbook(s): Tags: Body: LEM 9.6 was successfully run for the North Linac on Thu Nov 15 15:33:20 2018. The application phase took 0:01:35 to complete.

FSD Masking Tool: Successful System Masking for: BeamLossMonitor

Log Number: 3626718Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: lrichard
Areas: All

BeamLossMonitor Devices:
ILM1L02
ILM0I07
ILM6R08
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0R04
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R06
ILM0R08
ILM0R09
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T09
ILM4C01
ILM4C02
ILM4C03
ILM1C01
ILM1C

COMMENT ON *RF System Repair & Recovery - 2018

Log Number: 3626714Logbook(s): Body: 

2L13 CPS: CPS blower was making strange noises. FML was
notified and they replaced the belt. Blower noise went away
after replacement. Zone restored to High Voltage. Techs: D
Fasenbaker and staff.

Pages